# JUTRACK: a Julia package for auto-differentiable accelerator modeling and particle tracking

### v1

@@ Abstract
Efficient accelerator modeling and particle tracking are key for the design and configuration of modern particle accelerators. In this work, we present JUTRACK, a nested accelerator modeling package developed in Julia programing language and enhanced with compiler-level automatic differentiation (AD). With the aid of AD, JUTRACK enables rapid derivative calculations in accelerator modeling, facilitating sensitivity analyses and optimization tasks. We demonstrate the effectiveness of AD-derived derivatives through several practical applications, including sensitivity analysis of space-charge-induced emittance growth, nonlinear beam dynamics analysis for a synchrotron light source, and lattice parameter tuning of the future Electron-Ion Collider (EIC). Through the incorporation of automatic differentiation, this package opens up new possibilities for accelerator physicists in beam physics studies and accelerator design optimization.
@@ 1 Introduction
The design and optimization of modern particle accelerators require precise modeling of beam dynamics in complex electromagnetic environments, accounting for strong nonlinear interactions within the beam. Over the past few decades, many successful accelerator modeling tools have been developed, such as MAD-X [5], ELEGANT [6], AT [7], BMAD [8], IMPACT [9, 10] and TRACK [11]. These tools have been instrumental in advancing accelerator physics, providing reliable simulations that have significantly contributed to the design and optimization of numerous accelerator facilities worldwide. Recently, there has been growing interest in developing the digital twin of accelerator facilities [12]. The modeling tools will be an essential part of the digital twin for predicting the beam behaviors in the accelerator. Nevertheless, modern accelerators, such as the diffraction-limited storage ring (DLSR) [13], and future colliders, e.g., the Electron-Ion Collider (EIC) [14], the Future Circular Collider (FCC) [15], the circular electron positron collider (CEPC) [16], and muon colliders [17], have become increasingly complex and challenging in the dynamics of the beam. The highly nonlinear beam dynamics in the accelerators strongly limits the dynamic aperture and beam lifetime, making the optimization process extremely computationally intensive.
@@
Auto-differentiation (AD) has emerged as a powerful technique in computational physics and machine learning [19, 20, 21], offering notable benefits in sensitivity analysis and optimization tasks. AD enables automatic computation of derivatives, which are essential for gradient-based optimization methods and for analyzing the behavior of physical systems under small perturbations. Applying AD to accelerator modeling has significant potential to improve the efficiency of optimizing complex accelerator systems and provide deeper insights into intricate beam dynamics. In recent years, several new software packages that incorporate AD capabilities have been developed. One such innovative package is Cheetah [22], a PyTorch-based differentiable linear beam dynamics code. Another notable example is BMAD, a successful accelerator modeling tool in use since the mid 1990s, which is also planning to transition to Julia with AD capabilities to further enhance its computational performance [23].
@@
In this work, we present a new AD-enabled particle tracking package, JUTRACK. This package is written in Julia [24], a high-level, high-performance programming language specifically designed for scientific computing. With the aid of Enzyme [25], an AD plugin working at compiler level, JUTRACK allows automatic and rapid calculation of the derivatives of functions. In addition to standard particle tracking, this package also supports Truncated Power Series Algebra (TPSA) [26] tracking where the new particle coordinates are represented by a set of polynomials of initial coordinates. This package offers optional features for simulating collective effects, including transverse space-charge effects, wakefield effects, and beam-beam interactions.
@@ 2.1 Package overview
Figure 1 illustrates the functionalities of the package, which consists of two main components: the physics code and the TPSA functions. The primary goal of this package is to ensure that every target physics quantity such as beam emittance and every TPSA coefficient obtained from physics simulation are differentiable with respect to parameters of interest, which typically include the design parameters of the accelerator lattice and the initial coordinates of particles.
@@ 2.2 Implementation of automatic differentiation
To implement AD in JUTRACK, we utilize a third-party plugin, Enzyme [25]. Enzyme is a high-performance AD compiler plugin for the LLVM compiler framework, which is capable of calculating derivatives of statically analyzable programs expressed in the LLVM intermediate representation (IR). This compiler plugin is available in any language whose compiler targets LLVM IR, including C, C++, Fortran, Julia, Rust and others. Unlike conventional source-to-source and operator-overloading tools, Enzyme implements AD at the compiler level, resulting in significantly better computational efficiency [28].
@@ 2.3 Implementation of particle tracking
JUTRACK supports 6-D tracking of particle’s coordinates, $x$, $p_{x}$, $y$, $p_{y}$, $z$, and $\delta$. $x$ and $y$ are the transverse coordinates, $p_{x}$ and $p_{y}$ are the transverse momentum, $z=-\beta c\tau$ is the path lengthening with respect to the reference particle with $\tau$ as the relative arrival time at location $s$, and $\delta=dE/E_{0}$ is the energy deviation with respect to the reference particle.
@@ 2.4 Multivariate TPSA module
The TPSA, first introduced by Berz [26], has been used in acceleartor modeling code since 1990s [30]. A TPSA variable can be represented as a vector of Taylor series coefficients. The operations on the TPSA variables, such as addition and multiplication, are evaluated as vector operations following the rules of TPSA. The single-variable TPSA can be extended to the multivariate form [31]. Taking 6-D variables in our particle tracking regime as an example, the 6-D coordinates can be represented as,
@@
To implement multivariate TPSA, we have developed overloaded operators for a wide range of functions, including addition, subtraction, multiplication, division, trigonometric, inverse trigonometric, hyperbolic, square root, power, exponential, and logarithmic functions. These operators cover all types of operations required in our particle tracking code. By using these established TPSA variables and overloaded operators, standard particle tracking can be seamlessly replaced by tracking TPSA variables.
One important use of the TPSA module in JUTRACK is to calculate the $6\times 6$ linear transfer map $M$ of the accelerator. Considering the 6-D TPSA coordinates [$x$, $p_{x}$, $y$, $p_{y}$, $z$, $\delta$] at a starting location in the accelerator, we can calculate the coordinates after one turn. The first-order coefficients of these final coordinates form the linear transfer map $M$. This transfer matrix is fundamental for calculation of the optics functions, such as the $\beta$ functions and dispersion functions.
@@ 3.1 Validation of accuracy
To evaluate the correctness of particle tracking module in JUTRACK, the 6-D trajectory of an electron with random initial transverse coordinates [$x$, $y$] is tracked within the ESR. The results are compared with those of the well-benchmarked Polymorphic Tracking Code (PTC) integration of MAD-X [32], the software used for the EIC design. The comparison is illustrated in Fig. 2. The discrepancies in the final coordinates [$x$, $p_{x}$, $y$, $p_{y}$, $z$, $\delta$] between the two codes are close to the round-off error of the computer, which are negligible.
@@ 3.2 Performance evaluation
Without implementing parallel computing, the computational efficiency of multi-particle tracking is of the same order of magnitude for both packages. This implies Julia provides similar or even better efficiency compared to traditional low level languages. Additionally, JUTRACK’s parallel computing capability significantly reduces the computation time by an order of magnitude. The efficiency of JUTRACK can be further enhanced with more powerful computing devices equipped with additional processors, allowing for even larger and more complex simulations to be performed within practical time frames.
@@ 4.1 Sensitivity analysis of emittance growth due to space charge
where $K=qI/2\pi\epsilon_{0}p_{0}v_{0}^{2}\gamma_{0}^{2}$ is the generalized perveance, $a$ is the horizontal width of the pipe, $b$ is the vertical width of the pipe, $\alpha_{l}=l\pi/a$ and $\beta_{m}=m\pi/b$. The space charge potential $\phi^{lm}$ is given in the spectral domain as,
@@
where $N_{p}$ is number of particles and $\gamma^{2}_{lm}=\alpha_{l}^{2}+\beta_{m}^{2}$.
@@
In this example, we perform particle tracking simulations using 5000 macro particles and 12×12 spectral modes. The final emittance (normalized by the initial emittance) is calculated using JUTRACK particle tracking module. In the meantime, the derivatives of the final emittance with respect to the lattice parameters are automatically obtained via AD. Figure 4 presents these derivatives with respect to the lengths of the drift spaces (D1L, D2L and D3L), and the lengths (Q1L and Q2L) and strengths (Q1k and Q2k) of the quadrupoles. For validation purposes, we also compute approximate derivatives using the finite difference method. The close agreement between the AD results and the finite difference approximations confirms the accuracy of the AD computations. Our results show that the emittance variation is more sensitive to the length of the first quadrupole, which is consistent with the findings reported in [34].
@@ 4.2 Application to beam dynamics analysis
The dynamic aperture (DA) is a fundamental concept in the design and operation of storage ring accelerators. DA represents the volume in phase space where particles can circulate stably without being lost due to nonlinear effects. DA is usually determined through extensive particle tracking simulations, which involves tracking a large number of particles with varying initial conditions over a sufficiently long period in the accelerator. While accurate, this method is computationally expensive, often requiring significant computational resources and time. To fast optimize DA, accelerator physicists proposed an alternative approach, the resonance driving terms (RDTs) [35]. By calculating and minimizing these terms, one can enhance the DA, which is widely used in accelerator design and optimization.
@@
As a practical example, we apply JUTRACK to the Stanford Positron Electron Asymmetric Ring (SPEAR3) [37], a third-generation synchrotron light source at SLAC National Accelerator Laboratory. Five geometric RDTs, $h_{21000}$, $h_{10110}$, $h_{30000}$, $h_{10200}$, and $h_{10020}$, are calculated for the ring. To demonstrate the effectiveness of the AD-derived derivatives, we select two families of sextupoles installed in the SPEAR3 ring, namely, SDM and SFM, for the optimization of these RDTs. The derivatives of the RDTs with respect to the sextupole strengths can be automatically obtained in the calculation.
Using the derivatives, a simple gradient-descent optimization is implement to suppress the geometric RDTs. Figure 5 shows that the RDTs are reduced through the optimization. After approximately 500 iterations, the algorithm converged to a local minimum where the gradients with respect to both SDM and SFM sextupole strengths approached zero. This result indicates that the RDTs can be effectively minimized with JUTRACK, which could lead to an improvement in the DA. The calculation of the RDTs and their derivatives takes less than one second on a personal laptop, providing the potential for rapid optimization of the DA.
@@ 4.3 Application to design optimization
[…]
The rapid convergence to the optimal solution within just five steps highlights the remarkable efficiency and effectiveness in tuning the lattice parameters. This capability can significantly reduce the computational time and efforts typically required for lattice design optimization.
@@ Acknowledgement
This work is supported by DOE office of science, with award number DE-SC0024170.

### v2

@@ Abstract
Efficient accelerator modeling and particle tracking are key for the design and configuration of modern particle accelerators. In this work, we present JuTrack, a nested accelerator modeling package developed in the Julia programming language and enhanced with compiler-level automatic differentiation (AD). With the aid of AD, JuTrack enables rapid derivative calculations in accelerator modeling, facilitating sensitivity analyses and optimization tasks. We demonstrate the effectiveness of AD-derived derivatives through several practical applications, including sensitivity analysis of space-charge-induced emittance growth, nonlinear beam dynamics analysis for a synchrotron light source, and lattice parameter tuning of the future Electron-Ion Collider (EIC). Through the incorporation of automatic differentiation, this package opens up new possibilities for accelerator physicists in beam physics studies and accelerator design optimization.
@@ 1 Introduction
The design and optimization of modern particle accelerators require precise modeling of beam dynamics in complex electromagnetic environments, accounting for strong nonlinear interactions within the beam. Over the past few decades, many successful accelerator modeling tools have been developed, such as COSY INFINITY [5], MAD-X [6], ELEGANT [7], AT [8], BMAD [9], IMPACT [10, 11] and TRACK [12]. These tools have been instrumental in advancing accelerator physics, providing reliable simulations that have significantly contributed to the design and optimization of numerous accelerator facilities worldwide. Recently, there has been growing interest in developing the digital twin of accelerator facilities [13]. The modeling tools will be an essential part of the digital twin for predicting the beam behaviors in the accelerator. Nevertheless, modern accelerators, such as the diffraction-limited storage ring (DLSR) [14], and future colliders, e.g., the Electron-Ion Collider (EIC) [15], the Future Circular Collider (FCC) [16], the circular electron positron collider (CEPC) [17], and muon colliders [18], have become increasingly complex and challenging in the dynamics of the beam. The highly nonlinear beam dynamics in the accelerators strongly limits the dynamic aperture and beam lifetime, making the optimization process extremely computationally intensive.
@@
Auto-differentiation (AD) has emerged as a powerful technique in computational physics and machine learning [20, 21, 22], offering notable benefits in sensitivity analysis and optimization tasks. AD enables automatic computation of derivatives, which are essential for gradient-based optimization methods and for analyzing the behavior of physical systems under small perturbations. Applying AD to accelerator modeling has significant potential to improve the efficiency of optimizing complex accelerator systems and provide deeper insights into intricate beam dynamics. For instance, COSY INFINITY and MAD-X can provide derivatives in beam dynamics simulation with truncated power series algebra (TPSA) [23]. In recent years, several new software packages that incorporate AD capabilities have been developed. One such innovative package is Cheetah [24], a PyTorch-based differentiable linear beam dynamics code. Another notable example is BMAD, a successful accelerator modeling tool in use since the mid 1990s, which is also planning to transition to Julia with AD capabilities to further enhance its computational performance [25]. In addition, MAD-NG [26] incorporates generalized truncated power series algebra (GTPSA) [27], enabling it to compute derivatives of arbitrary order.
@@
In this work, we present a new AD-enabled particle tracking package, JuTrack. This package is written in Julia [28], a high-level, high-performance programming language specifically designed for scientific computing. With the aid of Enzyme [29], an AD plugin working at the compiler level, JuTrack allows automatic and rapid calculation of the derivatives of functions. In addition to standard particle tracking, this package also supports TPSA tracking, where the new particle coordinates are represented by a set of polynomials of initial coordinates. This package offers optional features for simulating collective effects, including transverse space-charge effects, wakefield effects, and beam-beam interactions.
@@ 2.1 Package overview
Figure 1 illustrates the functionalities of the package, which consists of two main components: the physics code and the TPSA functions. The primary goal of this package is to ensure that every target physics quantity such as beam emittance and every TPSA coefficient obtained from physics simulation is differentiable with respect to parameters of interest, which typically include the design parameters of the accelerator lattice and the initial coordinates of particles.
@@ 2.2 Implementation of automatic differentiation
There are several mature packages with AD capability that have been developed in the accelerator physics community, such as BMAD-X [31], Cheetah, and MAD-NG. BMAD-X is a Python package with library-agnostic tracking routines based on BMAD that allows high-dimensional gradient-based optimization, model calibration, and phase space reconstruction. Cheetah, built on PyTorch [32], integrates seamlessly with machine learning workflows, enabling advanced applications such as modular surrogate modeling and reinforcement learning. These Python-based packages are more accessible to a wider user base than traditional packages developed in non-scripting languages. Machine learning packges like Pytorch usually supports backward AD. It will be challenging to manage computing graphs for backward AD in complicated simulation code. MAD-NG, which utilizes GTPSA, provides flexibility with arbitrary-order differentiation. Considering in practical optimization tasks, people are usually more concerned with the first-order derivatives, we focus on the first-order derivative computation in the development of our package.
To implement AD to the first order with high efficiency, we utilize a third-party plugin, Enzyme [29]. Enzyme is a high-performance AD compiler plugin for the LLVM compiler framework with the capability of both forward and backward AD. It is capable of calculating derivatives of statically analyzable programs expressed in the LLVM intermediate representation (IR). This compiler plugin is available in any language whose compiler targets LLVM IR, including C, C++, Fortran, Julia, Rust, and others. Unlike conventional source-to-source and operator-overloading tools, Enzyme implements AD at the compiler level, resulting in significantly better computational efficiency [33]. Although Enzyme supports high-order derivatives computation, we only focus on first-order derivatives in this package for best efficiency. The users can extend it to higher-order if necessary.
@@ 2.3 Implementation of particle tracking
JuTrack supports 6-D tracking of particle’s coordinates, $x$, $p_{x}$, $y$, $p_{y}$, $z$, and $\delta$. $x$ and $y$ are the transverse coordinates, $p_{x}$ and $p_{y}$ are the transverse momentum, $z=-\beta c\tau$ is the path lengthening with respect to the reference particle with $\tau$ as the relative arrival time at location $s$, and $\delta=dp/p_{0}$ is the energy deviation with respect to the reference particle.
@@ 2.4 Multivariate TPSA module
The TPSA, first introduced by Berz [23], has been used in accelerator modeling code since 1990s [35]. A TPSA variable can be represented as a vector of Taylor series coefficients. The operations on the TPSA variables, such as addition and multiplication, are evaluated as vector operations following the rules of TPSA. The single-variable TPSA can be extended to the multivariate form [36]. Taking 6-D variables in our particle tracking regime as an example, the 6-D coordinates can be represented as,
@@
Multivariate TPSA is widely used for transfer matrix calculation and high-order transfer map analysis for beam dynamics. To implement multivariate TPSA, we have developed overloaded operators for a wide range of functions, including addition, subtraction, multiplication, division, trigonometric, inverse trigonometric, hyperbolic, square root, power, exponential, and logarithmic functions. These operators cover all types of operations required in our particle tracking code. By using these established TPSA variables and overloaded operators, standard particle tracking can be seamlessly replaced by tracking TPSA variables.
Considering the 6-D TPSA coordinates [$x$, $p_{x}$, $y$, $p_{y}$, $z$, $\delta$] at a starting location in the accelerator, we can calculate the coordinates after one turn. The first-order coefficients of these final coordinates form the linear transfer matrix, which is fundamental for the calculation of the optics functions, such as the $\beta$ functions and dispersion functions. In addition to TPSA, a finite difference approximation approach is also provided in this package for fast calculation of the first-order transfer map. The derivatives of the transfer map coefficients with respect to specific parameter, such as magnet strengths and initial coordinates of the particle, can be obtained with Enzyme.
@@ 3.1 Validation of accuracy
To evaluate the correctness of the particle tracking module in JuTrack, the 6-D trajectory of an electron with random initial transverse coordinates [$x$, $y$] is tracked within the ESR. The results are compared with those of the well-benchmarked Polymorphic Tracking Code (PTC) integration of MAD-X [37], the software used for the EIC design. The comparison is illustrated in Fig. 2. The discrepancy in the final coordinates [$x$, $p_{x}$, $y$, $p_{y}$, $z$, $\delta$] between the two codes is close to the round-off error of the computer, which is negligible.
@@ 3.2 Performance evaluation
Without implementing parallel computing, the computational efficiency of multi-particle tracking is of the same order of magnitude for both packages. This implies Julia provides similar or even better efficiency compared to traditional low-level languages. Additionally, JuTrack’s parallel computing capability significantly reduces the computation time by an order of magnitude. The efficiency of JuTrack can be further enhanced with more powerful computing devices equipped with additional processors, allowing for even larger and more complex simulations to be performed within practical time frames.
We also evaluate the efficiency of AD and compare it with the traditional numerical approach using finite differences. Our tests include a variety of functions, such as simple arithmetic operations with TPSA, particle tracking simulations, Twiss parameter calculations, and space charge computations. The computational times for these original functions ranged from microseconds to several minutes, depending on their complexity. For all tested functions, the computation time required for AD is almost twice that of the finite difference method, which is primarily due to the overhead associated with AD.
Despite the increased computational time, AD offers significant advantages in terms of accuracy and reliability. Finite difference methods are sensitive to the choice of step size and often struggle with numerical noise caused by truncation errors and floating point precision. In contrast, AD provides exact derivatives up to machine precision without the need to select an appropriate step size. This makes AD particularly advantageous for ill-conditioned functions, which are frequently seen in the nonlinear beam dynamics.
@@ 4.1 Sensitivity analysis of emittance growth due to space charge
where $K=qI/2\pi\epsilon_{0}p_{0}v_{0}^{2}\gamma_{0}^{2}$ is the generalized perveance, $\tau$ is the effective length, $i$ is the index of the $i$th particle, $a$ is the horizontal width of the pipe, $b$ is the vertical width of the pipe, $\alpha_{l}=l\pi/a$ and $\beta_{m}=m\pi/b$. The space charge potential $\phi^{lm}$ is given in the spectral domain as,
@@
where $N_{p}$ is the number of particles and $\gamma^{2}_{lm}=\alpha_{l}^{2}+\beta_{m}^{2}$ and $j$ is the index of the $j$th particle.
@@
In this example, we perform particle tracking simulations using 5000 macro particles and 12×12 spectral modes. The final emittance (normalized by the initial emittance) is calculated using the JuTrack particle tracking module. In the meantime, the derivatives of the final emittance with respect to the lattice parameters are automatically obtained via AD. Figure 4 presents these derivatives with respect to the lengths of the drift spaces (D1L, D2L and D3L), and the lengths (Q1L and Q2L) and strengths (Q1k and Q2k) of the quadrupoles. For validation purposes, we also compute approximate derivatives using the finite difference method. It is found that the results of finite difference approach nearly converge when the difference is less than 10-4. However, when the difference is too small, e.g., less than 10-9, the round-off errors will dominate. Therefore, we choose a reasonable difference of 10-6 in the calculation. The close agreement between the AD results and the finite difference approximations confirms the accuracy of the AD computations. Our results show that the emittance variation is more sensitive to the length of the first quadrupole, which is consistent with the findings reported in [39].
@@ 4.2 Application to beam dynamics analysis
The dynamic aperture (DA) is a fundamental concept in the design and operation of storage ring accelerators. DA represents the volume in phase space where particles can circulate stably without being lost due to nonlinear effects. DA is usually determined through extensive particle tracking simulations, which involve tracking a large number of particles with varying initial conditions over a sufficiently long period in the accelerator. While accurate, this method is computationally expensive, often requiring significant computational resources and time. To fast optimize DA, accelerator physicists proposed an alternative approach, the resonance driving terms (RDTs) [40]. By calculating and minimizing these terms, one can enhance the DA, which is widely used in accelerator design and optimization.
@@
As a practical example, we apply JuTrack to the Stanford Positron Electron Asymmetric Ring (SPEAR3) [42], a third-generation synchrotron light source at SLAC National Accelerator Laboratory. Five geometric RDTs, $h_{21000}$, $h_{10110}$, $h_{30000}$, $h_{10200}$, and $h_{10020}$, are calculated for the ring based on the lattice functions including Twiss parameters, dispersion, and phase advances. To demonstrate the effectiveness of the AD-derived derivatives, we select two families of sextupoles installed in the SPEAR3 ring, namely, SDM and SFM, for the optimization of these RDTs. The derivatives of the RDTs with respect to the sextupole strengths can be automatically obtained in the calculation.
Using the derivatives, a simple gradient-descent optimization is implemented to suppress the geometric RDTs. Figure 5 shows that the RDTs are reduced through the optimization. After approximately 500 iterations, the algorithm converged to a local minimum where the gradients with respect to both SDM and SFM sextupole strengths approached zero. This result indicates that the RDTs can be effectively minimized with JuTrack, which could lead to an improvement in the DA. The calculation of the RDTs and their derivatives takes less than one second on a personal laptop, providing the potential for rapid optimization of the DA.
@@ 4.3 Application to design optimization
With the AD capabilities of JuTrack, the derivatives of $\Delta\psi$ with respect to $k_{i}$ are calculated during the calculation of $\Delta\psi$. These derivatives are then utilized to perform a gradient-descent optimization, where the strengths $k_{1}$ to $k_{7}$ are simultaneously adjusted based on the derivatives with a fixed step size. For comparison, another optimization is conducted using derivatives approximated by the finite difference approach. Figure 6 illustrates the evolution of the multivariate optimization process. Within five iterations, the gradient-descent algorithm successfully converges to the optimal solution, achieving a phase advance $\Delta\psi$ of $2\pi$.
The results indicate that the derivatives obtained through AD and finite difference are almost the same. This is because the linear optics involved in the problem do not introduce complicated nonlinear correlations that could lead to ill-conditioned systems. As a result, the finite difference method is sufficiently accurate for this case. However, for more complicated problems, particularly those involving nonlinear forces from elements such as sextupoles, the finite difference approach may struggle due to its limited accuracy. In such scenarios, AD remains a reliable choice to provide precise derivatives.
@@ Acknowledgement
This work is supported by the DOE Office of Science, with award number DE-SC0024170.